An electron penetrates the region between two uniformly charged horizontal plates, but oppositely signed. Its initial speed, at half distance between the plates is 2×10 ^6 i m/s. The plates have a length of 4 cm and are 1.6 cm apart. In the whole region between the plates, the electric field ́ is uniform and perpendicular to the plates. (a) what is the maximum modulus that the vertical electric field can have so that the electron does not touch any plate?


The answer is 228 N/C but I'm not too sure how they got that?


1
Expert's answer
2022-02-16T13:24:17-0500

Velocity

"v=2\times10^6\hat{i}m/sec"

L=4cm

b=1.6cm

Time

"t=\frac{s}{v}"

"t=\frac{0.04}{2\times10^6}=2\times10^{-8}sec"

"s=ut+\frac{1}{2}at^2"

"0.005=0+0.5\times a\times4\times10^{-16}"


"a=\frac{0.005}{0.5\times4\times10^{-16}}=2.5\times10^{13}m/sec^2"

"a=\frac{eE}{m}"

"E=\frac{ma}{e}"

Put value


"E=\frac{9.1\times10^{-31}\times2.5\times10^{13}}{1.6\times10^{-19}}=228N/C"